How many rings are present in C16Hi3IO3? This compound consumes 3 mol of H on catalytic hydrogenation. Enter your answer in the provided box. ring(s)
Answer #1

Since 3 mol of H2 is required,

total number of double bonds = 3

Degree of unsaturation:

DOU = (2C + 2 + N - X - H)/2

X is number of halogens

C is number of carbon

N is number of Nitrogen

H is number of Hydrogen

use:

DOU = (2C + 2 + N - X - H)/2

= (2*16 + 2 + 0 - 1 - 23)/2

= (34 - 24) / 2

= 10/2

= 5

So,

total number of double bonds + total number of rings = 5

3 + total number of rings = 5

total number of rings = 2

Answer: 2